OpenStudy (anonymous):

Find the midpoint of the points (3, 1) and (7, –5

OpenStudy (anonymous):

take the average in each coordinate

OpenStudy (anonymous):

\(\frac{7+3}{2}=\frac{10}{2}=5\) and \(\frac{1-5}{2}=\frac{-4}{2}=-2\)

OpenStudy (anonymous):

your answer is therefore \((5,-2)\)